# A Modular Safety Filter for Safety-Certified Cyber-Physical Systems

### v1

@@ Abstract
Nowadays, many control systems are networked and embed communication and computation capabilities. Such control architectures are prone to cyber attacks on the cyberinfrastructure. Consequently, there is an impellent need to develop solutions to preserve the plant’s safety against potential attacks. To ensure safety, this paper introduces a modular safety filter approach that is effective for a variety of cyber-attack types. This solution can be implemented in combination with existing control and detection algorithms, effectively separating safety from performance. The safety filter does not require information on the reliability of the received command or the feature of the used anomaly detector. It can be implemented in conjunction with high-performance, resilient controllers, to achieve both high performance during normal operation and safety during an attack. As an illustrative example, we have shown the effectiveness of the proposed design considering a multi-agent formation task involving 20 mobile robots. The simulation results testify that the safety filter operates effectively during false data injection and intelligent attacks.
@@ I INTRODUCTION
Cyber-Physical Systems (CPS) are networked control systems with a tight integration with computation and communication capabilities[1]. Merging cyber technologies with physical systems greatly boosts operational efficiencies. However, it also introduces vulnerabilities that undermine the reliability of essential infrastructure as the communication lines present opportunities for hackers to manipulate data lines and initiate cyber attacks. A variety of solutions have been proposed to prevent, detect, and mitigate cyber-attacks using control theoretical tools[2]. Most of the solutions consider systems without constraints and/or the mitigation strategies rely on the use of an anomaly detector to ensure that the system does not enter unsafe configurations.
@@
Recently, there has been an increasing trend in constrained CPS to address safety concerns explicitly. This involves formally defining safe zones as constraints within state and input spaces. In this paper, a Modular Safety Filter (MSF), inspired by safety-certified learning-based controllers [3]-[4], is proposed to satisfy safety constraints in the presence of cyber-attacks on the actuator and sensor signals. Due to its modularity, this method can be used as a standalone technology alongside other resilient controllers and anomaly detectors. Performance and safety criteria are separated in our architecture, simplifying the design process. Since MSF makes no assumptions about the attacked signal, such as limited bandwidth and small bounds, or the attacker’s computational power, it can be used in a wide range of situations.
@@
Prior work largely focuses on linear systems. In [5] and [6], (set-theoretic) MPCs are employed for linear systems under False Data Injection (FDI) attacks, guaranteeing the stability and constraint satisfaction of the system. In [7], a distributed MPC and attack detection framework is proposed for constrained linear multi-agent systems under adversarial attacks. In [8], a method for tracking problems that require reachable sets is proposed for constrained linear systems under arbitrary attacks on both the actuation and measurement lines. A data-driven approach for LTI systems is proposed in [9] and [10], introducing a safety verification plus emergency control module assuming only noise-polluted input-state trajectories are available. A semi-definite approach, assuming bounded additive attacker’s signals, is proposed in [11] to design a safety-preserving filter for deterministic LTI systems under FDI attacks. In [12], a set-theoretic receding horizon control has been proposed to address FDI and denial of service attacks for LTI systems. In [13] and [14], reachability analysis is used and investigated to design safety-preserving platforms for LTI systems. A solution for nonlinear systems is proposed in [15] and [16] which provides safety based on an invariant set of SOS-based Lyapunov functions, resulting in conservative ellipsoidal safe sets.
The majority of proposed methods are developed resorting to reachable set arguments. This reliance often leads to a focus on LTI systems, since they become hard-to-compute or expensive for high-order and nonlinear systems. There is an emerging trend to separate the control performance from the safety. However, defining performance and safety through a single objective may degrade performance and result in conservative solutions where the safety objective is not directly linked to the control objective. To achieve superior performance, one may aim to optimize a non-continuous cost function via a deep policy while ensuring safety constraints in the presence of an attack, a task not feasible with current solutions. The proposed architecture aims to address these gaps.
Figure 1: Proposed safety-certified architecture for cyber-physical systems.
@@
The remainder of the paper is organized as follows. Preliminary material, adversarial capabilities, and the problem statement are described in section II. Section III includes the proposed MSF. A multi-agent setup consisting of 20 mobile robots is described in section IV. Simulation results obtained considering different FDI attacks are presented in section V. Lastly, conclusions and limitations of the present work are discussed in section VI.
II Problem Statement and Preliminary Material
@@
This section provides an overview of the CPS problem considered in this paper and the proposed MSF architecture, depicted in Fig. 1. It also details the formulation of the physical system and the types of cyber-attacks used for simulation examples.
II-A Physical System
@@
Consider a class of dynamical systems that can be described by a set of non-linear equations as follows,
$\dot{x}(t)=f(x(t),u(t)),$ (1)
where $t\in\mathbb{R}$ is the continuous time, $x=x(t)\in\mathbb{R}^{n}$ is the state vector, and $u=u(t)\in\mathbb{R}^{m}$ is the control input. For a unique solution, $f(.)$ is assumed to be continuously differentiable, and $u$ is a piecewise-continuous signal.
Definition 1 (Admissible Set).
An admissible set for the dynamical system (1) is defined as a subset of the state- and input-space as follows:
$\begin{split}u\in\mathcal{U}\subseteq\mathbb{R}^{m},x\in\mathcal{X}\subseteq\mathbb{R}^{n}.\end{split}$ (2)
Definition 2 (Safe Set).
A safe set, $\mathcal{S}\subseteq\mathcal{X}$, is a set of initial conditions for which there exists an input-state trajectory subjected to the dynamical system (1) that belongs to the admissible set (2) for an infinite time.
$\mathcal{S}=\{x(t_{0})\in\mathcal{X}\mid\exists u(t)\in\mathcal{U},x(t)\in\mathcal{X},\forall t>t_{0}\}.$ (3)
@@ II-B Adversarial Capabilities
Let us assume that $x(t)$ is sent from the local network to the cyber layer via an unsecured network. The cyber layer computes the control action, $u_{c}=u_{c}(t)$, and transmits it back to the local network. These signals are susceptible to cyber-attacks, denoted by $x_{a}=x_{a}(t)$ and $u_{a}=u_{a}(t)$, respectively. Note that this scenario is similar to the scenarios considered in [9] and [11]. Without losing generality, a cyber attack can be described using the following unknown function:
$(u_{a},x_{a})=h(u_{c},x).$ (4)
@@
Function $h(\cdot,\cdot)$ can be defined to represent different types of attack and adversarial capabilities [^1] The attacker may use previous inputs and states. Due to the simplicity of the notation, it is omitted here.. While the proposed modular safety filter does not rely on assumptions on the attack, i.e., the function $h(\cdot,\cdot)$ is unknown to the proposed safety filter, we will consider three situations in the illustrative simulation example in Section V: attack-free, FDI, and intelligent attacks.
II-B1 Attack-free
$h(\cdot,\cdot)$ is defined as the identity function in the absence of the attack as follows:
$u_{a}=u_{c},\ x_{a}=x.$ (5)
II-B2 FDI Attack
The attacker has the capability of adding arbitrary perturbations $\delta_{x}$ and $\delta_{u}$ to the control input and state measurement vectors.
$u_{a}=u_{c}+\delta_{u},\ x_{a}=x+\delta_{x}.$ (6)
II-B3 Intelligent Attack
The attacker is assumed to know the system dynamics and safe set, $\mathcal{S}$, as well as disclosure and disruptive resources on $x(t)$ and $u_{c}(t)$. Furthermore, the attacker has sufficient computational power to compute $h(\cdot,\cdot)$ resorting to any desired optimal policy. This includes an intelligent undetectable attack strategy that cannot be identified by an anomaly detector in the cyber layer. In general, a generic attack action can be modeled as the following optimization problem.
$\max\quad J(x,u_{c})$ (7)
@@ II-B3 Intelligent Attack
where $J(\cdot,\cdot)$ is the attacker’s objective and the attacker’s constraints represents conditions to ensure the attack is undetectable by an anomaly detector. Such conditions depend on the used anomaly detector and are left unspecified for the sake of generality.
II-C Central Controller
The central control policy may be designed for regulation, tracking, or any other task.
$u_{c}=g(r,x_{a}),$ (8)
where $r$ represents the reference trajectory. It is assumed that the controller is safety-certified in the absence of FDIs. The function $g(\cdot,\cdot)$ is unknown to the proposed safety filter [^2] The central controller may use future reference and past input-state data. Due to the simplicity of the notation, it is omitted here..
III Modular Safety Filter
We propose an architecture for which a modular safety filter can be implemented as an independent add-on module by filtering the control signal. Upon receiving $x(t)$ and $u_{a}(t)$, the safety filter provides the nearest safe input, $u_{s}(t)$, to the command signal $u_{a}(t)$, while respecting the system constraints (2) at all times. The safety filter can be described by the following optimization problem,
$u_{s}$ $:=\operatorname*{arg\,min}_{u}\|u_{a}-u\|^{2}_{2}$ (9)
s.t. safety $\quad\text{\emph{safety constraints}},\quad\forall t>0$
where ||.$||.||_{2}$ is the 2-norm of a vector. The safety constraints represents satisfying condition (2). To solve this problem, inspired by [17], we propose a predictive-based safety filter to handle safety constraints at all times including when potentially unsafe inputs are presented. The predictive safety filter approximates the problem (9) by searching for a safe backup trajectory toward a final safe set with finite-time prediction. Consider a discrete version of (1) as follows,
$x(k+1)=f_{d}(x(k),u_{s}(k)),$ (10)
where $u_{s}(k)=u_{0}^{k}$ is the first element of the solution of the optimization (III). The discrete-time index is defined as $k=\frac{t}{T_{s}}$ with $T_{s}$ representing the sampling time. The predictive safety filter is outlined below,
$\min_{u_{n\in\mathcal{N}}^{k}}\|u_{a}(k)-u_{0}^{k}\|^{2}_{2}$ (11a)
[…]
$x_{n+1}^{k}=f_{d}(x_{n}^{k},u_{n}^{k}),$ (11c)
$(x_{n}^{k},u_{n}^{k})\in(\mathcal{X},\mathcal{U}),$ (11d)
$x_{N}^{k}\in\mathcal{S}_{f},$ (11e)
$x_{0}^{k}=x(k),$ (11f)
where (11a) yields the nearest safe action to $u_{a}(k)$, (11c) is the prediction model, (11d) is the admissible set, (11e) is the final safe set, (11f) is the initial condition at discrete time $k$, $u_{n}^{k}$ is the $n^{th}$ element of prediction at discrete-time index $k$, and $N$ is the prediction horizon. Note that if $u_{a}$ is safe, the safety filter does not alter the input. The solution to this optimization problem yields an input-state backup trajectory. The safe set $\mathcal{S}_{f}$ is defined to guarantee the recursive feasibility of the filter similar to the terminal condition in MPC [18]. A visual illustration of the admissible set, safe set, final safe set, and the backup trajectory at time $k$ is shown in Fig. 2. The safety filter algorithm is summarized in the Algorithm 1.
@@ III Modular Safety Filter
(Initial Feasibility) The optimization problem (III) is feasible at $k=0$.
Assumption 2.
(Final Safe Set, $\mathcal{S}_{f}$) Let $\mathcal{S}_{f}\subseteq\mathcal{X}$, and $\mathcal{K}_{f}$ be a local policy that renders system (10) invariant on $\mathcal{S}_{f}$.
Note that the safe set, $\mathcal{S}$, is implicitly considered via the MSF optimization problem (III). The size of this set depends on the prediction horizon and the size of the final safe set, $\mathcal{S}_{f}$. A short prediction horizon and a small final safe set may lead to an overly conservative solution. Generally, there is no need for a long prediction horizon to achieve a non-conservative solution, in contrast to a solution that aims to provide a robust performance controller as well as safety.
@@
Proof.
The proof follows from Theorem 4.6 in [17]. ∎
Note that safety remains ensured if the (non-convex) optimization is infeasible for some $k>0$, as a safe backup trajectory is available [17].
Figure 2: At the time $k$, an unsafe control input, $u_{a}(k)$, is received by the safety filter. Since applying this input may result in an unsafe trajectory (I) in the next steps, MSF will find a backup trajectory (III) towards the final safe set, $\mathcal{S}_{f}$, by applying the safe input, $u_{s}(k)$. Applying this safe input results in the modified trajectory (II).
Algorithm 1 Predictive Safety Filter for CPS
1:Initialize $\mathcal{S}_{f}$, $\mathcal{X}$, $\mathcal{U}$, $N$, $x(0)$, $k=0$.
@@
3:     Solve problem (III) for $u_{a}(k)$.
4:     Apply $u_{s}(k)$ to system (1).
5:     Measure system’s states, $x(k+1)$, send it to the cyber layer and update the initial condition, $x_{0}^{k+1}$.
6:     $k\xrightarrow{}k+1$
@@
Remark: In this paper, we assumed that the underlying problem is deterministic, nominal, and has zero transmission delay. For other settings when a simplified model, probabilistic model, or additive disturbance is present, an adaptation of the algorithm is required which may introduce conservatism, see [3] and[4] for a recent overview of safety filter technology.
IV Centralized Predictive Safety Filter for Multi-Agent Mobile Robot System
To evaluate the efficiency of the proposed method on a high-order nonlinear system, we adopt the simulation framework employed in [19], which considered 20 linear mobile robots, extended with nonlinear dynamics as described in [20]. For simplicity, we assume that all mobile robots have the same dynamics and parameters. The kinematics model for the $i^{th}\in\mathcal{I}$ robot is described as follows:

### v2

@@ Abstract
Nowadays, many control systems are networked and embed communication and computation capabilities. Such control architectures are prone to cyber attacks on the cyberinfrastructure. Consequently, there is an impellent need to develop solutions to preserve the plant’s safety against potential attacks. To ensure safety, this paper introduces a modular safety filter approach that is effective for various cyber-attack types. This solution can be implemented in combination with existing control and detection algorithms, effectively separating safety from performance. The safety filter does not require information on the received command’s reliability or the anomaly detector’s feature. It can be implemented in conjunction with high-performance, resilient controllers to achieve both high performance during normal operation and safety during an attack. As an illustrative example, we have shown the effectiveness of the proposed design considering a multi-agent formation task involving 20 mobile robots. The simulation results testify that the safety filter operates effectively during undetectable, intelligent attacks.
@@ I INTRODUCTION
Cyber-Physical Systems (CPS) are networked control systems with a tight integration with computation and communication capabilities[1]. Merging cyber technologies with physical systems significantly boosts operational efficiencies. However, it also introduces vulnerabilities that undermine the reliability of essential infrastructure as the communication lines present opportunities for hackers to manipulate data lines and initiate cyber attacks. Various solutions have been proposed to prevent, detect, and mitigate cyber-attacks using control theoretical tools[2]. Most of the solutions consider systems without constraints, and/or the mitigation strategies rely on the use of an anomaly detector to ensure that the system does not enter unsafe configurations.
@@
Recently, there has been an increasing trend in constrained CPS to address safety concerns explicitly. This involves formally defining safe zones as constraints within state and input spaces. In this paper, a Modular Safety Filter (MSF), inspired by safety-certified learning-based controllers [3]-[4], is proposed to satisfy safety constraints in the presence of cyber-attacks on the actuator and sensor signals. Due to its modularity, this method can be used as a standalone technology alongside other resilient controllers and anomaly detectors. Performance and safety criteria are separated in our architecture, simplifying the design process. Since MSF makes no assumptions about the attacked signal, such as limited bandwidth and small bounds or the attacker’s computational power, it can be used in a wide range of situations.
@@
Prior work largely focuses on linear systems. In [5] and [6], Model Predictive Control (MPC) is employed for linear systems under False Data Injection (FDI) attacks, guaranteeing the stability and constraint satisfaction of the system. In [7], a distributed MPC and attack detection framework is proposed for constrained linear multi-agent systems under adversarial attacks. In [8], a tracking method that requires reachable sets is proposed for constrained linear systems under arbitrary attacks on both the actuation and measurement lines. A data-driven approach for LTI systems is proposed in [9] and [10], introducing a safety verification plus emergency control module, assuming only noise-polluted input-state trajectories are available. A semi-definite approach, assuming bounded additive attacker’s signals, is proposed in [11] to design a safety-preserving filter for deterministic LTI systems under FDI attacks. In [12], a set-theoretic receding horizon control has been proposed to address FDI and denial of service attacks for LTI systems. In [13] and [14], reachability analysis is used and investigated to design safety-preserving platforms for LTI systems. A solution for nonlinear systems is proposed in [15] and [16], which provides safety based on an invariant set of SOS-based Lyapunov functions, resulting in conservative ellipsoidal safe sets.
The majority of proposed methods rely on reachable set arguments, which often restrict their focus to LTI systems, as computing reachable sets becomes challenging or costly for high-order and nonlinear systems. Additionally, many methods aim to address both control performance and safety, requiring a balance between these objectives and computational cost — a challenge for MPC, particularly with long horizons in high-order systems. In contrast, a modular safety filter approach ensures safety without modifying the existing system or requiring long prediction horizons, making it computationally efficient and practical for nonlinear systems. This paper explores how CPS safety can be maintained without altering existing components, such as controllers and anomaly detectors, by incorporating a minimally invasive filter. MSF facilitates the integration of established control methods, achieving both performance and safety without system modifications.
Figure 1: Proposed safety-certified architecture for cyber-physical systems. The green dashed lines (Refer to caption) represent the communication between the modular safety filter and the plant, assumed to be unaffected by network attacks. The attacker can target communication channels between the plant and the networked controller, represented by gray solid lines (Refer to caption), and injects malicious signals through red dotted lines (Refer to caption). As an illustrative example, this architecture can be applied to a robotic scenario, where the networked controller communicates with the robot via Wi-Fi, with the onboard modular safety filter operating, leaving the Wi-Fi connection vulnerable to potential threats. See the following color box for a detailed discussion of the proposed architecture.
[!t] Why safety filters must be local to the plant. It is important to emphasize that a covert, intelligent attack on communication channels may evade detection by any anomaly detectors on the networked controller’s side [17]. This underscores the need for a safety filter to act as a localized policy on the plant side. In CPS domains like the smart grid, the networked controller oversees and synchronizes subsystems to achieve a unified goal while managing only each subsystem’s reference/input signal; see discussions in [18] and [11]. Therefore, the safety filter requires access to the unaltered states, which can only be achieved through a localized implementation. However, in cases where the local network faces a sophisticated, intelligent attack, modifications to the filter may be necessary to maintain safety [19]. Finally, suppose an attacker executes a coordinated, intelligent attack. In that case, the only effective defense for the system is to implement an emergency safety module, such as the modular safety filter, local to the plant.
@@
The remainder of the paper is organized as follows. Preliminary material, adversarial capabilities, and the problem statement are described in section II. Section III includes the proposed MSF. A multi-agent setup consisting of 20 mobile robots is described in section IV. Simulation results considering undetectable, intelligent attacks are presented in Section V. Lastly, conclusions and limitations of the present work are discussed in section VI.
II Preliminary Material and Problem Statement
@@
This section provides an overview of the CPS problem considered in this paper, the proposed MSF architecture, and the networked controller depicted in Fig. 1. It also details the formulation of the plant’s dynamics and the types of cyber-attacks used for simulation examples.
II-A Plant’s Dynamics
@@
Consider a class of discrete-time dynamical systems that can be described by a set of nonlinear equations as follows,
$x(t+1)=f(x(t),u(t)),$ (1a)
$u(t)\in\mathcal{U},\;x(t)\in\mathcal{X},$ (1b)
where $t\in\mathbb{N}$ is the time step, $x(t)\in\mathbb{R}^{n}$ the state, $u(t)\in\mathbb{R}^{m}$ the control input, $\mathcal{U}$ the input constraints, and $\mathcal{X}$ the state constraints. The function $f(x(t),u(t))$ is a generic function describing the plant’s dynamic.
Definition 1 (Safety).
The dynamical system (1) is said to be safe if the input-state pair $(u(t),x(t))$ satisfies $(u(t),x(t))\in\mathcal{U}\times\mathcal{X}$ for all $t\geq 0$.
Definition 2 (Safe Control Invariant Set).
A safe control invariant set, $\mathcal{S}\subseteq\mathcal{X}$, is a set of initial states $x(t)$ such that there exists a control input $u(t)\in\mathcal{U}$ ensuring that $x(t+1)\in\mathcal{S}$ for all $t>0$. Formally:
$\mathcal{S}=\{x(t)\in\mathcal{X}\mid\exists u(t)\in\mathcal{U},\ x(t+1)\in\mathcal{S},\forall t>0\}.$ (2)
@@ II-B Adversarial Capabilities
Let us assume that $x(t)$ is sent from a local network to the networked controller via an unsecured network. The networked controller computes the control action, $u_{c}(t)$, and transmits it back to the local network. These signals are susceptible to cyber-attacks, denoted by $x_{a}(t)$ and $u_{a}(t)$, respectively. Without losing generality, a cyber attack can be described using the following unknown function:
$(u_{a}(t),x_{a}(t))=h(u_{c}(t),x(t)).$ (3)
@@
Function $h(u_{c}(t),x(t))$ can be defined to represent different types of attack and adversarial capabilities. Since the proposed modular safety filter does not rely on assumptions on the attack, i.e., the function $h(u_{c}(t),x(t))$ is unknown to the proposed safety filter. We will consider two situations in the illustrative simulation example in Section V: attack-free and intelligent attacks.
II-B1 Attack-Free Scenario
In the attack-free scenario, $h(u_{c}(t),x(t))$ is the identity function, indicating that the attacker cannot alter the signals $x(t)$ and $u_{c}(t)$:
$u_{a}(t)=u_{c}(t),\quad x_{a}(t)=x(t),$ (4)
In other words, the networked controller receives unaltered $x(t)$ from the local network, and the local network receives unaltered $u_{c}(t)$ from the networked controller.
II-B2 Intelligent Attack
The attacker is assumed to know the system dynamics, disclosure, and disruptive resources on the data transmitted, $x(t)$ and $u_{c}(t)$; undetectable covert attacks can be launched. Furthermore, the attacker has sufficient computational power to compute $h(u_{c}(t),x(t))$ resorting to any desired optimal policy. For this attack, it has been proved that no anomaly detector - whether implemented as an active or passive module - located on the networked controller’s side can detect its presence [20]. In particular, for an FDI attack, we assume that the attacker can introduce arbitrary perturbations, $\delta_{x}(t)$ and $\delta_{u}(t)$, to the control input and state measurement vectors:
$u_{a}(t)=u_{c}(t)+\delta_{u}(t),\quad x_{a}(t)=x(t)+\delta_{x}(t).$ (5)
II-C Networked Controller
A networked controller typically consists of two main components: a tracking controller and an anomaly detector. The tracking control policy can also be formulated to address tasks such as regulation, tracking, or other objectives. It is generally expressed as:
$u_{c}(t)=g(r(t),x_{a}(t)),$ (6)
where $r(t)$ represents the reference trajectory. The tracking controller is assumed to be safety-certified under the attack-free condition in (4). Note that violating this assumption does not compromise the plant’s safety, but it may trigger false alarms in anomaly detection systems (see Remark 2). Therefore, the function $g(r(t),x_{a}(t))$ is unknown to the proposed safety filter [^1] The attacker may exploit previously recorded inputs and states, as seen in replay buffer attacks, while the tracking controller may use future reference and past input-state data. For simplicity, this notation is omitted here.. We consider a passive binary anomaly detection mechanism designed to detect cyber-attacks. The anomaly detector is described as follows:
$a(t)=\mathcal{A}(\{x_{a}(i)\}_{i=t_{0}}^{t},\{u_{c}(i)\}_{i=t_{0}}^{t}),$ (7)
where $a(t)\in\{0,1\}$, with $a=0$ and $a=1$ indicating the absence and presence of an anomaly, respectively. Here, $\mathcal{A}(\{x_{a}(i)\}_{i=t_{0}}^{t},\{u_{c}(i)\}_{i=t_{0}}^{t})$ is a generic function that can store an arbitrarily large history of the signals $x_{a}(i)$ and $u_{c}(i)$, where $0\leq t_{0}\leq t$. It is important to note that in the case of an intelligent attack, no anomaly detection mechanism can reliably detect the anomaly on the networked controller side. Therefore, the anomaly detector in (7) is only included for completeness and is intended for use in simulation examples.
Problem 1.
Design a local control policy to ensure that system (1) remains safe, as defined in Definition 1, regardless of the attacker strategy (3). The policy should also preserve the functionality of the tracking controller (6) and anomaly detector (7) under attack-free conditions (4).
III Modular Safety Filter
To solve Problem 1, we propose an architecture for which a modular safety filter can be implemented as an independent add-on module by filtering the control signal. Upon receiving $x(t)$ and $u_{a}(t)$, the safety filter provides the nearest safe input, $u_{s}(t)$, to the command signal $u_{a}(t)$, while respecting the system constraints (1b) for all $t\geq 0$. The safety filter can be described by the following optimization problem,
$u_{s}(t)$ $=\operatorname*{arg\,min}_{u}\|u_{a}(t)-u(t)\|^{2}_{2}$ (8)
s.t. $\quad u(t)\in\mathcal{U},\quad x(t)\in\mathcal{X},\quad\forall t,\geq 0$
where ||.$||.||_{2}$ is the 2-norm of a vector. To solve this problem, inspired by [21], we propose a predictive-based safety filter to handle safety constraints at all times, including when potentially unsafe inputs are presented. The predictive safety filter approximates the problem (8) by searching for a backup input-state trajectory toward a terminal safe control invariant set with finite-time prediction. The predictive safety filter is outlined below,
$u_{s}(t)=\operatorname*{arg\,min}_{u_{t}^{k}}\|u_{a}(t)-u_{t}^{k}\|^{2}_{2}$ (9a)
@@ III Modular Safety Filter
$x_{t}^{k+1}=f(x_{t}^{k},u_{t}^{k}),$ (9c)
$(x_{t}^{k},u_{t}^{k})\in(\mathcal{X},\mathcal{U}),$ (9d)
$x_{t}^{N}\in\mathcal{S}_{f},$ (9e)
$x_{t}^{0}=x(t),$ (9f)
where (9a) yields the nearest safe action to $u_{a}(t)$, (9c) is the prediction model, (9d) is the admissible set, (9e) is the terminal safe control invariant set , (9f) is the initial condition at time $t$, $u_{t}^{k}$ is the $k^{th}$ element of prediction at time $t$, and $N$ is the prediction horizon. Note that if $u_{a}(t)$ respects constraints (9c)-(9f), the safety filter does not alter the input. The solution to this optimization problem yields an input-state backup trajectory at time step $t$ as $(u_{t}^{k},x_{t}^{k})$ for $k\in\mathcal{N}$. The safe set $\mathcal{S}_{f}$ is a control invariant set and defined to guarantee the recursive feasibility of the filter similar to the terminal condition in MPC [22], e.g., the equilibrium of point of the system (1). A visual illustration of the state constraints, safe set, terminal safe control invariant set, and the backup trajectory at time $t$ is shown in Fig. 2. The safety filter algorithm is summarized in the Algorithm 1.
@@
(Initial Feasibility) The optimization problem (9) is feasible at $k=0$.
Assumption 2 (Terminal Safe Control Invariant Set , $\mathcal{S}_{f}$).
The terminal safe control invariant set $\mathcal{S}_{f}$ is a known safe control invariant set satisfying $\mathcal{S}_{f}\subseteq\mathcal{S}$ under the terminal control policy $u_{\mathcal{S}_{f}}=K_{\mathcal{S}_{f}}(x)$.
Note that the safe set, $\mathcal{S}$, is implicitly considered via the MSF optimization problem (9). The size of this set depends on the prediction horizon and the size of the terminal safe control invariant set, $\mathcal{S}_{f}$. Generally, there is no need for a long prediction horizon to achieve a non-conservative solution when only safety is concerned, in contrast to an MPC solution that aims to provide performance and safety.
@@
Proof. To ensure safety, it is sufficient to prove that the optimization (9) enjoys recursive feasibility. If the optimization (9) admits a solution at $t$, then it means that there exists a sequence of control input $\left\{u_{t}^{0},\ldots,u_{t}^{N-1}\right\}$ that takes the state trajectory safely inside the control invariant region $\mathcal{S}_{f}$. Consequently, at $t+1$, one admissible although not optimal solution for (9) always exists, and it is given by $\left\{u_{t}^{1},\ldots,u_{t}^{N-1},u_{\mathcal{S}_{f}}.\right\}$. This is sufficient to ensure the recursive feasibility of (9) and, consequently, the existence of a safe backup trajectory provided by (9) regardless of the attacker’s actions. $\square$
Figure 2: At the time $k$, an unsafe control input, $u_{a}(k)$, is received by the safety filter. Since applying this input may result in an unsafe trajectory (I) in the next steps, MSF will find a backup trajectory (III) towards the terminal safe control invariant set, $\mathcal{S}_{f}$, by applying the safe input, $u_{s}(k)$. Applying this safe input results in the modified trajectory (II).
Algorithm 1 Modular Predictive Safety Filter for CPS
1:Initialize $\mathcal{S}_{f}$, $\mathcal{X}$, $\mathcal{U}$, $N$, $x(0)$, $t=0$.
@@
3:     Solve problem (9) for $u_{a}(t)$.
4:     Apply $u_{s}(t)$ to system (1).
5:     Measure system’s states, $x(t+1)$, send it to the networked controller and update the initial condition.
6:     $t\xrightarrow{}t+1$
@@
Remark 1: In this paper, we assumed that the underlying problem is deterministic, nominal, and has zero transmission delay. For other settings, when a simplified model, probabilistic model, or additive disturbance is present, an adaptation of the algorithm is required, which may introduce conservatism; see [3] and[4] for a recent overview of safety filter technology.
Remark 2: It is assumed that the tracking controller does not activate the safety filter in the absence of an attack, ensuring that MSF does not interfere with the anomaly detector unless an attack occurs on the communication channel. To relax this assumption, an additional copy of the MSF can be placed alongside the tracking controller. In this configuration, a pre-filtered control signal is transmitted, while the MSF on the plant side remains inactive in the absence of an attack [23].
IV Modular Predictive Safety Filter for a Multi-Agent Mobile Robot System
To evaluate the efficiency of the proposed method on a high-order nonlinear system, we adopt the simulation framework employed in [24], which considered 20 mobile robots. For simplicity, we assume that all mobile robots have the same dynamics and parameters. Unlike [24], which employs linear models, we utilize a nonlinear kinematics model for the $i^{th}$ robot, where $i\in\mathcal{I}$, described as follows: